OpenStudy (anonymous):

find the equivalent expression using the same bases. (21x28)^2 explain.

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

is that 21 times 28?

OpenStudy (anonymous):

i tried solving it but you dont multiply?

OpenStudy (anonymous):

its suppose to be something like the product rule or quotient rule idk

OpenStudy (anonymous):

have you tried multiplying 21 and 28 then applying the exponent on the product?

OpenStudy (anonymous):

yeah and it said express your answer in terms of 21 and 28 only

OpenStudy (anonymous):

oh you just have to distribute the exponents. like for example (3 x 4)^2 = 3^2 x 4^2 so if you'll apply it to your problem, that would be?

OpenStudy (anonymous):

hmm i think let me try the example said a^m x a^n=a ^m+n or a^m/a^n= a^m-n or i think yeah this one (a x b)^n= a^nxb^n yeah this is the step for it

OpenStudy (anonymous):

yes, so your answer will be.............?

OpenStudy (anonymous):

euh let me solve it first and then see if its right

OpenStudy (anonymous):

so i set it up like 21x28^2= 21^2x28^2 so that would equal 588^4 idk i messed up i think

OpenStudy (anonymous):

\(21\times 28^2= 21^2 \times 28^2\) is right. I think this is enough you dont have to evaluate it because they are asking only "in terms of 28 and 21"

OpenStudy (anonymous):

ohh right , i always end up solving lol that wasnt that hard, maybe my brain is finally understanding lol

OpenStudy (anonymous):

what about 39 x 96^8 cause it doesnt have two exponents and when i add them its number is way too big hmm...

OpenStudy (anonymous):

oh nvm i got it :)

OpenStudy (anonymous):

still need any help? sorry i was away again lol

OpenStudy (anonymous):

ill tag you in my new question lol and its fine :)